# Stochastic Thermodynamics of Micromagnetics with Spin Torque

### v1

@@ I Introduction
In the present work, we consider a micro-magnet driven by a non-conservative torque called spin-current torque or spin-torque, which arises due to interaction with a spin-polarized current. Although interaction between spin-current and magnetization is of quantum origin Stiles2006 , the resulting dynamics of the magnetization can be still described by the classical LL equation augmented by a new term Slonczewski1996 ; Berger1996 ; Brataas-Nature-materials-2012 . In this work, we shall incorporate stochasticity into this dynamics and study its stochastic thermodynamics. We shall find a novel effect called entropy pumping, which we argue represents the exchange of entropy between the system and the spin-current that does not involve dissipation. The effect of entropy pumping was firstly discovered by Kim and Qian Kim2004 ; Kim2007 some time ago in Hamiltonian system driven by velocity dependent forces.
This work is organized as follows. In Sec. II we derive the stochastic Landau-Lifshitz-Slonczewski (s-LLS) equation, which describes the Langevin dynamics of a micro-magnet coupled both to a magnetic field and to a spin-torque, and discuss the conditions of local detailed balance. In Sec. III we develop the theory of stochastic thermodynamics. We establish the first and second laws of thermodynamics, and discuss the effect of entropy pumping. In Sec. IV we prove a fluctuation theorem for the joint distribution of work and pumped entropy, and verify it using numerical simulation. Finally in Sec. V we draw the concluding remarks.
@@ II The Stochastic Landau-Lifshitz-Slonczewski equation
where the vector $\bm{P}$ is parallel to the spin-polarization of the current, which in general may depend on many details of the spin-current, as well as on the magnetization ${\bm{m}}$ itself. We shall make the simplifying assumption that $\bm{P}$ is independent of ${\bm{m}}$. Below we shall call $\bm{P}$ the spin current vector, for simplicity. Hence ${\bm{m}}\times\bm{P}$ may be understood as the nonconservative effective magnetic field induced by the spin-polarized current. It is nonconservative because if cannot be expressed as the gradient of certain potential function. The resulting equation shall be called stochastic Landau-Lifshitz-Slonczewski (s-LLS) equation :
@@
We may decompose the current into a reversible current $j^{i}_{\rm R}$ and an irreversible current $j^{i}_{\rm IR}$:
@@
In the absence of spin torque, Eq. (15) reduces to Eq. (2.20) of Ref. LLG-1 , and Eqs. (16) reduce to Eqs. (2.21) of Ref. LLG-1 .
@@ II.1 Detailed balance
For micromagnetic systems without spin torque, the conditions of detailed balance were discussed in Sec. IID of Ref. LLG-1 . These conditions guarantee that the steady state can be understood as a thermodynamic equilibrium. In the presence of spin torque, these conditions should be properly generalized, and should be called the conditions of local detailed balance. The conditions of local detailed balance for the covariant Langevin equation (7) driven by non-conservative forces were given in Eqs. (3.4) and (3.6) of Ref. sto-therm-NC . In the present case, the magnetization ${\bm{m}}$ plays the role of state variable, and the magnetic field $\bm{H}$ plays the role of control parameter. Note that both ${\bm{m}}$ and $\bm{H}$ are odd under time-reversal, whereas the spin current vector $\bm{P}$ is even under time-reversal, which means that $\varepsilon_{i}=-1$, $\bm{H}^{*}=-\bm{H},\bm{P}^{*}=\bm{P}$. (The fact that $\bm{P}$ is even under time-reversal can be easily seen via dimensional analysis of Eq. (5).) Hence Eqs. (3.4) and (3.6) of Ref. sto-therm-NC become
@@ III Stochastic thermodynamics
A general theory of stochastic thermodynamics was developed for non-conservative Langevin dynamics (7) in Ref. sto-therm-NC . This theory may be directly applied to s-LLS equation (6).
@@
To appreciate the physical meanings of work defined above, we consider the special case of vanishing damping coefficient $\eta=0$, the s-LLS equation (6) reduces to
@@
${dS^{\rm tot}}={dS}+{dS_{\rm env}}.$ (30)
Whilst $dS$ is the differential of Gibbs-Shannon entropy:
@@
${dS_{\rm env}}$ should be understood as the ensemble average of Eq. (29). After some tedious calculation, we obtain (c.f. Eq. (3.78) of Ref. sto-therm-NC ):
@@ IV Fluctuation theorem
Taking advantage of Eq. (48), and using standard methods of stochastic thermodynamics, we can prove the following fluctuation theorem for $\mathscr{W}$ and $\mathscr{S}_{\rm P}$:
@@
which may be understood as the stochastic entropy production along $\bm{\gamma}$, if the final state of the system is an equilibrium state. We can then use Eq. (49) to prove the following fluctuation theorem:

### v2

@@ I Introduction
In the present work, we consider a micro-magnet driven by a non-conservative torque called spin-current torque or spin-torque, which arises due to interaction with a spin-polarized current. Although interaction between spin-current and magnetization is of quantum origin Stiles2006 , the resulting dynamics of the magnetization can be still described by the classical Landau-Lifshitz equation augmented by a new term Slonczewski1996 ; Berger1996 ; Brataas-Nature-materials-2012 . In this work, we shall incorporate stochasticity into this dynamics and study its stochastic thermodynamics. We shall find a novel effect called entropy pumping, which we argue represents the exchange of entropy between the system and the spin-current that does not involve dissipation. The effect of entropy pumping was firstly discovered by Kim and Qian Kim2004 ; Kim2007 some time ago in Hamiltonian system driven by velocity dependent forces.
This work is organized as follows. In Sec. II we derive the stochastic Landau-Lifshitz-Slonczewski (sLLS) equation, which describes the Langevin dynamics of a micro-magnet coupled both to a magnetic field and to a spin-torque, and discuss the conditions of local detailed balance. In Sec. III we develop the theory of stochastic thermodynamics. We establish the first and second laws of thermodynamics, and discuss the effect of entropy pumping. In Sec. IV we derive the fluctuation theorem for the joint distribution of work and pumped entropy, and verify it using numerical simulation. Finally in Sec. V we draw the concluding remarks.
@@ II The Stochastic Landau-Lifshitz-Slonczewski equation
where the vector $\bm{P}$ is parallel to the spin-polarization of the current, which in general may depend on many details of the spin-current, as well as on the magnetization ${\bm{m}}$ itself. We shall make the simplifying assumption that $\bm{P}$ is independent of ${\bm{m}}$. Below we shall call $\bm{P}$ the spin current vector, for simplicity. Hence ${\bm{m}}\times\bm{P}$ may be understood as the nonconservative effective magnetic field induced by the spin-polarized current. It is nonconservative because if cannot be expressed as the gradient of certain potential function. The resulting equation shall be called stochastic Landau-Lifshitz-Slonczewski (sLLS) equation :
@@
We may decompose the current into a reversible current $j^{i}_{\rm R}$ and an irreversible current $j^{i}_{\rm IR}$ sto-therm-NC :
@@
These names are pertinent because, as we will show below in Eq. (33), the irreversible current $j^{k}_{\rm IR}$ but not the reversible current $j^{k}_{\rm R}$ contributes to the entropy production. Invoking Eqs. (3d) and (3e), we see that the reversible current $j^{k}_{\rm R}$ is proportional to the gyromagnetic ratio $\gamma_{0}$, whereas the irreversible current $j^{k}_{\rm IR}$ is proportional to the damping coefficient $\eta$. We note that in the absence of spin torque, Eq. (15) reduces to Eq. (2.20) of Ref. LLG-1 , and Eqs. (16) reduce to Eqs. (2.21) of Ref. LLG-1 .
@@ II.1 Detailed balance
For micromagnetic systems without spin torque, the conditions of detailed balance were discussed in Sec. IID of Ref. LLG-1 . These conditions guarantee that the steady state can be understood as a thermodynamic equilibrium. In the presence of spin torque, these conditions should be properly generalized, and should be called the conditions of local detailed balance. The conditions of local detailed balance for the covariant Langevin equation (7) driven by non-conservative forces were given in Eqs. (3.4) and (3.6) of Ref. sto-therm-NC . In the present case, the magnetization ${\bm{m}}$ plays the role of state variable, and the magnetic field $\bm{H}$ plays the role of control parameter. Note that both ${\bm{m}}$ and $\bm{H}$ are odd under time-reversal, whereas the spin current vector $\bm{P}$ is even under time-reversal. (The fact that $\bm{P}$ is even under time-reversal can be easily seen via dimensional analysis of Eq. (5).) Hence Eqs. (3.4) and (3.6) of Ref. sto-therm-NC become
@@ III Stochastic thermodynamics
A general theory of stochastic thermodynamics was developed for non-conservative Langevin dynamics (7) in Ref. sto-therm-NC . This theory may be directly applied to sLLS equation (6).
@@
To appreciate the physical meanings of work defined above, we consider the special case of vanishing damping coefficient $\eta=0$, the sLLS equation (6) reduces to
@@
[…]
where $dS$ is the differential of Gibbs-Shannon entropy:
@@
whilst ${dS_{\rm env}}$ is the ensemble average of Eq. (27). After some tedious calculation, we obtain (c.f. Eq. (3.78) of Ref. sto-therm-NC ):
@@ IV Fluctuation theorem
Taking advantage of Eq. (48), and using standard methods of stochastic thermodynamics, we can establish the following fluctuation theorem for $\mathscr{W}$ and $\mathscr{S}_{\rm P}$:
@@
which may be understood as the stochastic entropy production along $\bm{\gamma}$, if the final state of the system is an equilibrium state. We can then use Eq. (49) to establish the following fluctuation theorem: